OpenStudy (anonymous):

find the equation for the sequence: 2.1, 2.02, 2.003, 2.0004

Directrix (directrix):

Is this a finite sequence?

Directrix (directrix):

Interesting that this sequence comes up on a Crack site.

OpenStudy (anonymous):

2.1 = 2 + 1/10 2.02 = 2 + 2/100 2.003 = 2 + 3/1000 Does that help enough?